OpenStudy (anonymous):

Is the following problem simplified correctly? (5x+7)(5x-7)=25x2+70x-49

myininaya (myininaya):

(5x+7)(5x-7)= 25x^2-49

myininaya (myininaya):

whenever you have (a-b)(a+b)=a^2-b^2

myininaya (myininaya):

\[5x(5x-7)+7(5x-7)=25x^2-35x+35x-49=25x^2-49\]